Write down the negation of the following statements and determine the truth value of the negation:

a) βˆ€π‘₯ ∈ 𝑹, π‘₯ 2 + 1 β‰₯ 2π‘₯

b) βˆ€π‘₯ ∈ 𝑹, (𝑦 β‰ β†’ (𝑦 + 1)/𝑦 < 1

c) βˆƒπ‘§ ∈ 𝒁, (𝑧 𝑖𝑠 π‘œπ‘‘π‘‘) ∨ (𝑧 𝑖𝑠 𝑒𝑣𝑒𝑛)

d) βˆƒπ‘› ∈ 𝑡, (𝑛 𝑖𝑠 𝑒𝑣𝑒𝑛) ∧ (βˆšπ‘› 𝑖𝑠 π‘π‘Ÿπ‘–π‘šπ‘’)


1
Expert's answer
2021-11-16T11:08:59-0500

a)"\exist x \isin R, x^2 +1<2x". False, because least value of "x^2 -2x +1=(x-1)^2" is 0 in case x=1.

b)"\exist y \isin R, y ≠ 0 ∧ (y+1)/y ≥ 1". True, because for example with any "y>0" , "1+1\/y \u22651".

c)"\forall z \isin Z," "(z \space is \space not \space odd) ∧ (z \space is \space not \space even)". False, because if we choose any odd or even integer, and statement become false.

d)"\forall n \isin N, (n \space is \space not \space even) ∨ (√n \space is \space not \space prime)". False, because for example n=4, in this case n is even and √n=√4=2 is prime.